# GABInsight: Exploring Gender-Activity Binding Bias in Vision-Language Models

### v1

@@ 1 Introduction
As far as we are aware, this research is the first to delve into the origins of the association between genders and activities in vision-language models by conducting a thorough analysis both image-to-text and text-to-image retrieval accuracy of VLMs. It encompasses not only straightforward scenarios involving single individuals, where VLMs demonstrate high accuracy for both genders [14, 3, 35], but also more complex situations with individuals of different genders [10]. This comprehensive approach allows us to isolate the impact of the complexity of the scenario on performance decline [38] and more accurately measure the Gender-Activity Binding bias. Additionally, the study carries out supplementary experiments to investigate the bias in VLMs’ text encoders and their proficiency in recognizing activities.
@@ 4 Dataset
Pipeline of the dataset creation is shown in Figure 1(a)
@@ 4.1 Selection of Activities
We proposed three methods to identify the biased activities, all utilizing Masked Language Modeling (MLM) as part of their approach to detecting biased activities. We utilized the ’RoBERTa-large’ [23] model for its superior performance in masked token prediction, which was attributed to its objective function. The sentences are formatted as "a <mask> is doing <activity>".444<mask> is the special mask token of the RoBERTa model We then calculate the probabilities associated with "man" and "woman" (as well as related masculine and feminine identifiers like "boy" and "girl"), from which we calculate the log ratios, as specified in Section 4.1.1.
@@ 4.1.1 Stereotyped Activities from GPT Combined with MLM
We applied this methodology separately for activities stereotypically associated with each gender. For men, this resulted in a normal distribution555all the tests including log ratios formed normal distributions that we tested using Shapiro–Wilk test for normality with a positive mean ($\mu>0$), from which we chose the sentences at the rightmost end. Similarly, for women, it also resulted in a normal distribution but with a negative mean ($\mu<0$), from which we then chose the sentences at the leftmost end.
@@ 5.2.3 Text-to-Image Retrieval
Figure 4 illustrates the accuracy of VLMs in the task of text-to-image retrieval. For each activity, we form random pairs of images from E2 and U2777Pairing is in the form of one-to-one correspondence.. Each model is then evaluated for its ability to retrieve the image that best matches captions formatted as "a <man/woman> is <doing activity> and a <woman/man> is in the scene". As can be seen in Figure 4, VLMs achieve an accuracy of approximately 50%, indicating that their performance is nearly random in this task. If we replicate this experiment, altering only the captions to a gender-neutral phrase "a person is <doing activity>", and designate the image where the expected gender is performing the activity as the true label, the VLMs achieve an accuracy of approximately 50% again, as illustrated in Figure 4 in the Appendix. In essence, the benchmarked models do not seem to comprehend the properties of the given images that would aid in retrieving the correct one given embedded information from the caption, i.e., they fail to recognize the performer of the activity based on the text.

### v3

@@ 1 Introduction
As far as we are aware, this research is the first to delve into the origins of the association between genders and activities in vision-language models by conducting a thorough analysis both image-to-text and text-to-image retrieval accuracy of VLMs. It encompasses not only straightforward scenarios involving single individuals, where VLMs demonstrate high accuracy for both genders [14, 3, 35], but also more complex situations with individuals of different genders [10]. This comprehensive approach allows us to isolate the impact of the complexity of the scenario on performance decline [38] and more accurately measure the Gender-Activity Binding bias. Additionally, the study carries out supplementary experiments to investigate the bias in VLMs’ text encoders and their proficiency in recognizing activities. [^4] The dataset and codes are available on Github: https://github.com/sharif-ml-lab/GABInsight
@@ 4 Dataset
Pipeline of the dataset creation is shown in Figure 1(a).
@@ 4.1 Selection of Activities
We proposed three methods to identify the biased activities, all utilizing Masked Language Modeling (MLM) as part of their approach to detecting biased activities. We utilized the ’RoBERTa-large’ [23] model for its superior performance in masked token prediction, which was attributed to its objective function. The sentences are formatted as "a <mask> is doing <activity>".555<mask> is the special mask token of the RoBERTa model We then calculate the probabilities associated with "man" and "woman" (as well as related masculine and feminine identifiers like "boy" and "girl"), from which we calculate the log ratios, as specified in Section 4.1.1.
@@ 4.1.1 Stereotyped Activities from GPT Combined with MLM
We applied this methodology separately for activities stereotypically associated with each gender. For men, this resulted in a normal distribution666all the tests including log ratios formed normal distributions that we tested using Shapiro–Wilk test for normality with a positive mean ($\mu>0$), from which we chose the sentences at the rightmost end. Similarly, for women, it also resulted in a normal distribution but with a negative mean ($\mu<0$), from which we then chose the sentences at the leftmost end.
@@ 5.2.3 Text-to-Image Retrieval
Figure 4 illustrates the accuracy of VLMs in the task of text-to-image retrieval. For each activity, we form random pairs of images from E2 and U2888Pairing is in the form of one-to-one correspondence.. Each model is then evaluated for its ability to retrieve the image that best matches captions formatted as "a <man/woman> is <doing activity> and a <woman/man> is in the scene". As can be seen in Figure 4, VLMs achieve an accuracy of approximately 50%, indicating that their performance is nearly random in this task. If we replicate this experiment, altering only the captions to a gender-neutral phrase "a person is <doing activity>", and designate the image where the expected gender is performing the activity as the true label, the VLMs achieve an accuracy of approximately 50% again, as illustrated in Figure 4 in the Appendix. In essence, the benchmarked models do not seem to comprehend the properties of the given images that would aid in retrieving the correct one given embedded information from the caption, i.e., they fail to recognize the performer of the activity based on the text.